One hose can fill a pool in 15 hours. A second can fill the pool in 10 hours. How long will it take to fill the pool using both hoses?
Answer by richwmiller(17219)

1/15+1/10=1/x
2/30+3/30=1/x
5/30=1/x
1/6=1/x
x=6